OpenStudy (sneville):

Rewrite the expression in terms of the given function or functions. (sec x + csc x) (sin x + cos x) - 2 - cot x; tan x A. 0 B. 2tan x C. tan x D. 2 + tan x

OpenStudy (sshayer):

\[\left( \sec x+\csc x \right)\left( \sin x+\cos x \right)=\left( \frac{ 1 }{ \cos x }+\frac{ 1 }{ \sin x } \right) \left( \sin x+\cos x \right)\] \[=\left( \frac{ \sin x+\cos x }{ \cos x \sin x} \right)\left( \sin x+\cos x \right)=\frac{ \left( \sin x+\cos x \right)^2 }{ \cos x \sin x }=?\]